(5.2.23) Assume that when adults with smartphones are randomly​ selected, 43​% use them in meetings or classes. If 10 adult smartphone users are randomly​ selected, find the probability that at least 8 of them use their smartphones in meetings or classes.
The probability is?
1
Expert's answer
2020-12-01T06:51:00-0500

p = 43% = 0.43

n = 10

Binomial probability:

"P(X=x)=C(n,x)\cdot p^x\cdot (1-p)^{n-x}"

"P(X=8)=\frac{10!}{8!2!}\cdot 0.43^8\cdot 0.57^2=0.0171"

"P(X=9)=\frac{10!}{9!1!}\cdot 0.43^9\cdot 0.57^1=0.0029"

"P(X=10)=\frac{10!}{10!0!}\cdot 0.43^10\cdot 0.57^0=0.0002"

"P(X\geq8)=P(X=8)+P(X=9)+P(X=10)="

"=0.0171+0.0029+0.0002=0.0202"


Answer: 0.0202